OpenStudy (anonymous):

the second term of geometric sequence is 3/4 and its fourth term is 3 what is the first term? and show the solution. please answer... thank you :))

OpenStudy (welshfella):

4th term = a1*r^3 and 2nd term = a1*r so 4th / 2nd = a1*r^3 / a1*r = r^2 r^2 = 3 / (3/4) = 3 *4 / 3 = 4 r = 2 take 2nd term:- a1*r = a1*2 = 3/4 so first term a1 = 3/4 / 2 = 3/8

OpenStudy (anonymous):

okay.. thanks :)